### v2

# Dynamics of states of infinite quantum systems as a cornerstone of the second law of thermodynamics

Abstract
We improve on our version of the second law of thermodynamics as a deterministic theorem for quantum spin systems [Wre22] in two basic aspects. The first concerns the general statement of the second law: spontaneous changes in an adiabatically closed system will always be in the direction of increasing (mean) entropy, which rises to a maximal value. In order to arrive at this statement, we have to extend the previous definition of adiabatic transformation, in order to include sudden interactions, and thereby generalize the “barrier model”. The cornerstone of the second law is then seen to be the fact that the dynamics may induce a basic structural transition between states of infinite systems in the limit of large times, which, physically, represent times much larger than specific relaxation times. Two specific examples concern the transition from pure to mixed states in two different universality classes of dynamics in one dimension, one being the exponential model, which has exactly soluble dynamics and does not display a phase transition, the other the Dyson model(s), for which neither the dynamics nor the statistical mechanics are exactly soluble, and which do exhibit a (ferromagnetic) phase transition, first demonstrated by Dyson. It is also shown, as a consequence of results of Albert and Kiessling on the Cloitre function, that there is strong graphical evidence that the dynamics of the Dyson models are chaotic for large times, and, thereby, that the mechanisms of approach to equilibrium differ for these two universality classes in a profound way, which depend exclusively on the dynamics, and not on the states and observables.
1 Introduction: states and observables for classical systems, chaotic states, and the smoothing property of densities
In his basic work [Di30], Dirac emphasized that the basic building blocks of quantum mechanics are states and observables, states being vectors in a (finite-dimensional) Hilbert space, and observables, self-adjoint operators acting on these states. In the present paper we shall see that it is imperative, in order to arrive at a precise statement of the second law of thermodynamics in the sense of Clausius, to generalize this notion of state to infinite dimensions, or, more precisely, to infinite number of degrees of freedom (ndof). In order to sharpen our intuition, it appears to be useful to undertake a preliminary investigation of the situation in connection with classical dynamical systems, which leads us to the probabilistic properties of (deterministic) dynamical systems [LM85]. Since generic dynamical systems are chaotic, we choose to look at one of the simplest chaotic systems, the so-called dyadic map [LM85].
Define $T_{2}$ by $T_{2}:X\to Y$ , where $X=[0,1]$ and $Y=[0,1]$, by
$T_{2}x=fr(2x)\equiv 2x\mbox{mod}1$ (1)
Above, $fr(x)\equiv(x-[x])$, where $[x]$ denotes the largest integer which is smaller or equal to x. Note that this mapping is not one-to-one (it is a so-called endomorphism). In fact, we see that the inverse image of a point $x$ is either $x/2$ or $\frac{x+1}{2}$. $T_{2}$ is called the dyadic transformation and leaves Lebesgue measure $\mu$ (on the line) invariant because the inverse image of a point $y$ is
$T_{2}^{-1}(y)=\{\frac{y}{2}\}\cup\{\frac{y+1}{2}\}$ (2)
Indeed, from (2), $\mu(T_{2}^{-1}([0,y]))=y$, which generalizes to
$\mu(A)=\mu(T_{2}^{-1}(A))\mbox{ for any Borel set }A\in[0,1]$ (3)
$T_{2}$ has a simple alternative description. Let $x$ be given by its expansion in basis 2, i.e.,
$x=.\epsilon_{1}\epsilon_{2}\cdots=\sum_{n=1}^{\infty}\frac{\epsilon_{n}}{2^{n}}\mbox{ with }\epsilon_{n}\in\{0,1\}$ (4)
Then, it is easy to see that
$T_{2}x=.\epsilon_{2}\epsilon_{3}\cdots\mbox{ and, in general }T_{2}^{n}x=.\epsilon_{n+1}\epsilon_{n+2}\cdots$ (5)
that is, $T_{2}$ is the “one-sided shift” in this representation. Given a Lebesgue integrable function, i.e.,
$f\in L^{1}(0,1)$ (6)
we may define the Ruelle-Perron-Frobenius operator $P$ ([LM85]) by
$\int_{A}(Pf)(x)dx=\int_{T_{2}^{-1}(A)}f(x)dx$ (7)
If
$f_{0}(x)\equiv 1$ (8)
it follows that
$Pf_{0}\equiv 1$ (9)
Let $A=[0,x]$. From (7), it follows that
$(Pf)(x)=\frac{d}{dx}(\int_{0}^{\frac{x}{2}}f(u)du+\int_{\frac{x}{2}}^{\frac{x+1}{2}}f(u)du=$
$=\frac{1}{2}[f(\frac{x}{2})+f(\frac{x+1}{2})]$
from which, by iteration,
$(P^{n}f)(x)=\frac{1}{2^{n}}\sum_{k=0}^{2^{n}-1}f(\frac{x+k}{2^{n}})$ (10)
We shall say that f is a density if $f\geq 0$ almost everywhere in $(0,1)$ $\int_{0}^{1}f(x)dx=1$; $f_{0}$, given by (8), is the uniform density. The words a.e. stand for almost everywhere, that is, in the complement of a set with (Lebesgue) measure zero. By (7), P maps densities to densities, and, by (10),
$\lim_{n\to\infty}(P^{n}(f))(x)=\int_{0}^{1}f(y)dy=1$ (11)
Together with (8), (11) is a form of the property of approach to equilibrium (see also [LM85], Theorem 4.4.1, p.65): the evolution of any density tends (for (discrete) time $n\to\infty$) to the unique invariant density, the uniform density $f_{0}$ given by (8).
Of particular importance above is that $P$ is not defined on individual orbits ot the map, which would correspond to taking $f$ in (7) to be a “delta function”, which is not allowed by condition (6). Indeed, these individual orbits behave rather erratically. Consider two points $x_{1},x_{2}$, both in [0, 1], close in the sense that the first n digits in the expansion (9) are identical: it follows that $T_{2}x_{1}$ and $T_{2}x_{2}$ differ already in the first digit: an initially exponentially small difference $2^{-n}$ is magnified by the evolution to one of order O(1). When this property holds for $n$ arbitrarily large, as in the present example, one speaks of the exponential sensitivity to initial conditions.
Exponential sensitivity to initial conditions is a precise definition of the chaotic element of the dynamics: for almost all $x_{0}$ (in the sense of Lebesgue measure, i.e., excluding a set of zero Lebesgue measure) $\gamma$, $T_{2}^{n}x_{0}$ comes arbitrarily close to almost any $x\in[0,1]$, if $n$ is taken sufficiently large, or, in other words, it “fills” the whole interval uniformly throughout the evolution. Therefore $\lim_{n\to\infty}T_{2}^{n}x_{0}$ does not exist for a.e. $x_{0}$. The set $\gamma$ consists of the finite dyadic numbers $x_{f}$, i.e., those whose dyadic expansion  (4) is finite; it is immediate from the definition that $T_{2}^{n}x_{f}\to 0\mbox{ or }1$, the latter being the fixed points of the dyadic map: they are untypical in the sense that they do not fill the interval uniformly.
We now come to the subject of states and observables for the dyadic map and chaotic states. For an alternative treatment of states of classical systems, see Ruelle’s article [Ru67]. See also the paper by Illner and Neunzehrt [IN87], in which the change between the two topologies as below is held responsible for the fact that in the limit of a process with time-reversal invariance in the Boltzmann equation a “time-arrow” is generated. In order to characterize these two topologies, we begin with a definition.
Definition 1.1.
A density $f$ is a function $f\geq 0$ a.e. in $(0,1)$ such that $\int_{0}^{1}f(x)dx=1$. $f_{0}$, given by  (8), is the uniform density.
The present model may strengthen our intuition behind our choices of states and observables. Indeed, taking $f$ as a bounded continuous function on $[0,1]$, as an observable, and, as a state, the delta measure
$\omega_{1}(f)\equiv\delta_{x}(f)\mbox{ for }x\in[0,1]$ (12)
then the “time-evolution” leads to
$\omega_{1,n}(f)=\delta_{x_{n}}(f)=f(x_{n})$ (13)
It follows, however, that, for a.e. $x_{0}$, $x_{n}=T_{2}^{n}x_{0}$ comes arbitrarily close to almost any $x\in[0,1]$ [LM85]. Thus, $f(x_{n})$ cannot converge (for a.e. $x_{0}$), and there is no approach to equilibrium.
Consider, now, as observables, the densities defined in  (1.1). Define, now, as observable the entropy $\eta$ by $\eta(u)\equiv-u\log(u)\mbox{ with }\eta(0)\equiv 0$, and consider the states
$\omega_{2}(f)\equiv\int_{0}^{1}dx\eta(f)(x)$ (14)
By  (14), together with Theorem 9.2.1 and remark 9.2.1 of [LM85], the sequence of “time-evolved states”, where $P$ denotes the Ruelle-Perron-Frobenius operator  (7),
$\omega_{2,n}(f)\equiv\int_{0}^{1}dx\eta(P^{n}(f))(x)$ (15)
with $n=1.2,\cdots$, is monotone increasing in $n$, and bounded above by zero, and therefore converges. Thus: there is approach to equilibrium! This is the smoothing property of densities. It is due to the fact that densities require, for their definition, a (non-countably) infinite set of points, i.e., they must be supported on a set of non-zero Lebesgue measure. Equation  (15) is the analogue of the forthcoming definition of the mean entropy as a functional of states of infinite systems in quantum statistical mechanics.
We may finally observe, for the purpose of further comparison with the quantum case, that the state $\omega_{2}$ above is manifestly not invariant under “time-reversal”, since the dynamics, specified by $P$, is not. In fact, definitions  (14) of $\omega_{2}$ and  (7) of $P$ show that, for suitable “initial” observables $f$, $\omega_{2}(f)$ exhibits exponential sensitivity to initial conditions, due to the fact that the dynamics $T_{2}$ has this property: we say that $\omega_{2}(f)$ is a chaotic state.
The present model seems thus to suggest that the main issue in the problem of irreversibility and approach to equilibrium in quantum mechanics may also lie in the choice of suitable states and observables, in particular their non-invariance under time-reversal. We now examine this issue in greater detail.
2 Dynamics of states and the Schrödinger paradox. The case of states of infinite systems
In the present year, which marks the centennary of (Schrödinger) quantum mechanics, it may be specially significant to handle a topic to which Schrödinger provided, in 1935, the initial crucial input: the Schrödinger paradox [Schr65]. This topic is the theory of irreversibility and the second law of thermodynamics, already treated in [Wre22], but in an incomplete way. In this paper we attempt to supply the missing steps and, at the same time, clarify the principles underlying the choice between the von Neumann and the Boltzmann entropy, which, to our knowledge, has not been done before.
2.1 Models of qss and their dynamical properties
A typical example of the models we shall use is the generalized Heisenberg Hamiltonian (generalized Ising model (gIm) if $J_{1}=0$):
$H_{\Lambda}=-2\sum_{x,y\in\Lambda}[J_{1}(x-y)(S_{x}^{1}S_{y}^{1}+S_{x}^{2}S_{y}^{2})+J_{2}(x-y)S_{x}^{3}S_{y}^{3}]$ (16)
where
$\sum_{x\in\mathbf{Z}^{\nu}}|J_{i}(x)|<\infty\mbox{ and }J_{i}(0)=0\mbox{ for }i=1,2$ (17)
[…]
The notion of state generalizes to systems with infinite number of degrees of freedom $\omega(A)=\lim_{\Lambda\nearrow\infty}\omega_{\Lambda}(A)$, at first for $A\in{\cal A}_{L}$ and then to $\cal A$}.
Each state $\omega$ defines a representation $\Pi_{\omega}$ of $\cal A$} as bounded operators on a Hilbert space ${\cal H}_{\omega}$ with cyclic vector $\Omega_{\omega}$ (i.e., $\Pi_{\omega}({\cal A})\Omega_{\omega}$ is dense in ${\cal H}_{\omega}$), such that $\omega(A)=(\Omega_{\omega},\Pi_{\omega}(A)\Omega_{\omega})$ (the GNS construction). The strong closure of $\Pi_{\omega}({\cal A})$ is a von Neumann algebra, with commutant $\Pi_{\omega}({\cal A})^{{}^{\prime}}$, which is the set of bounded operators on ${\cal H}_{\omega}$ which commute with all $\Pi_{\omega}({\cal A})$, and the center is defined by $Z_{\omega}=\Pi_{\omega}({\cal A})\cap\Pi_{\omega}({\cal A})^{{}^{\prime}}$.
Considering quantum spin systems on $\mathbf{Z}^{\nu}$, we shall consider only space-translation-invariant states, i.e., such that
$\omega\circ\tau_{x}=\omega\mbox{ for all }x\in\mathbf{Z}^{\nu}$ (18)
For states of finite ndof, we assume periodic boundary conditions.
A pure, or ergodic state is a state which cannot be written as a proper convex combination of two distinct states $\omega_{1}$ and $\omega_{2}$, i.e., the following does not hold:
$\omega=\alpha\omega_{1}+(1-\alpha)\omega_{2}\mbox{ with }0<\alpha<1$ (19)
If the above formula is true, it is natural to regard $\omega$ as a mixture of two pure “phases” $\omega_{1}$ and $\omega_{2}$, with proportions $\alpha$ and $1-\alpha$, respectively.
Let $\Gamma=\mathbf{Z}^{\nu}$ and $P_{0}(\Gamma)$ denote the collection of all finite parts of $\Gamma$. We consider in this paper quantum spin systems described by a Hamiltonian for any finite region $\Lambda\subset\mathbf{Z}^{\nu}$
$H_{\Phi}(\Lambda)=\sum_{X\subset\Lambda}\Phi(X)$ (20)
where $\Phi$ is a translation-invariant interaction function from $P_{0}(\Gamma)$ into the Hermitian elements of the quasi-local algebra $\cal A$}, such that $\Phi(X)\in{\cal A}(X)$ and
$\Phi(X+x)=\tau_{x}(\Phi(X))\forall X\subset\mathbf{Z}^{\nu},\forall x\in\mathbf{Z}^{\nu}$ (21)
We shall for simplicity restrict ourselves to quantum spin systems with two-body interactions. In this case, stability requires that
[…]
A prototype of the above is the anisotropic ferromagnetic Heisenberg model: it corresponds to the choices:
[…]
where $n.n.$ denote nearest-neighbors in the lattice $\mathbf{Z}^{\nu}$. If $J_{2}\equiv 0$, we speak of the X-Y model. The assumption $J_{1}\equiv 0$ in  (16), with $J_{2}=J$ satisfying  (17), defines the generalized Ising model (gIm) studied by Emch [Em66] and Radin [Ra70]:
$H_{\Lambda}=\frac{1}{2}\sum_{x,y\in\Lambda}J(|x-y|)\sigma_{x}^{3}\sigma_{y}^{3}$ (24)
Let, now, $\nu=1$, and define two subclasses of models of the gIm:
$J(|x|)=-\xi^{-|x|}\mbox{ with }\xi>1\mbox{ and }x\neq 0(\mbox{ \emph{exponential model} }E_{\xi})$ (25)
and
[…]
The above conditions on $\xi$ and $\alpha$ are required for stability (20). Under certain conditions, the Dyson model displays a ferromagnetic phase transition [Dy69]. With $P_{0}(\mathbf{Z}^{\nu})$ denoting the set of all finite subsets of $\mathbf{Z}^{\nu}$ as before, for each triple $A=(A_{1},A_{2},A_{3})$, where the $A_{i}\in P_{0}(\mathbf{Z}^{\nu},i=1,2,3$ are pairwise disjoint, define
$\sigma^{A}\equiv\prod_{x_{1}\in A_{1}}\sigma_{x_{1}}^{1}\prod_{x_{2}\in A_{2}}\sigma_{x_{2}}^{2}\prod_{x_{3}\in A_{3}}\sigma_{x_{3}}^{3}$ (27)
Let $\omega$ be any state satisfying
$\omega(\sigma^{A})=0\forall A\mbox{ such that }A_{3}\neq\phi$ (28)
We have the following:
Theorem 2.1.
For the gIm with interaction either of the exponential model $E_{\xi}$ with $\xi$ a transcendental number, or the Dyson model $D_{\alpha}$, and any initial state (at $t=0$) $\omega$ satisfying (28),
$\lim_{t\to\infty}(\omega\circ\tau_{t})=\omega_{eq}\equiv\otimes_{x\in\mathbf{Z}}(2^{-1}Tr_{x})$ (29)
where the limit on the l.h.s. of (29) is taken in the weak* topology.
The equilibrium state above is the tracial state, defined, for general dimension $\nu$, as
$\omega_{tr}\equiv\otimes_{x\in\mathbf{Z}^{\nu}}(2^{-1}Tr_{x})$ (30)
where the infinite sum over $x$ is short for the limit in the weak*-topology of the set of sums over finite regions.
For the proof, see ([Ra70], Corollary, p.2953).
Quantum spin systems are similar to quantum fields because of the now famous Lieb-Robinson bound ([LR72], see also [BRo297], p.254). Following the last reference, let us assume, instead of (20),
[…]
for some $\lambda>0$. Then, for $A,B\in{\cal A}_{0}$,
[…]
Equation (32) means that physical disturbances propagate with “group velocity” bounded by
[…]
2.2 The mean entropy, the Schrödinger paradox and the time-arrow: the necessity of states of infinite systems
For a finite quantum spin system the von Neumann entropy is defined as ($k_{B}=1$):
$S_{\Lambda}=-Tr(\rho_{\Lambda}\log\rho_{\Lambda})$ (34)
We may view $\rho_{\Lambda}$ as a state $\omega_{\Lambda}$ on ${\cal A}(\Lambda)$ - a positive, normed linear functional on ${\cal A}(\Lambda)$: $\omega_{\Lambda}(A)=Tr_{{\cal H}_{\Lambda}}(\rho_{\Lambda}A)\mbox{ for }A\in{\cal A}(\Lambda)$ (positive means $\omega_{\Lambda}(A^{{\dagger}}A)\geq 0$, normed $\omega_{\Lambda}(\mathbf{1})=1$.)
We may correspondingly view $S_{\Lambda}$ as a functional of the state $\omega_{\Lambda}$. Accordingly, $S_{\Lambda}(t)$ will be a functional $S_{\Lambda}(\omega_{\Lambda,t})$ of the time-evolved state
$\omega_{\Lambda,t}\equiv Tr_{{\cal H}_{\Lambda}}(\rho_{\Lambda}\tau_{t}^{\Lambda}(A))$ (35)
where
$\tau_{t}^{\Lambda}(A)\equiv\exp(iH_{\Lambda}t)A\exp(-iH_{\Lambda}t)\forall A\in{\cal A}(\Lambda)$ (36)
By  (34) and  (35) and the cyclicity and invariance of the trace under unitary transformations, it follows that
$\omega_{\Lambda,t}=\omega_{\Lambda,-t}$ (37)
We now come back to the second law. Assume, as we shall always do in this paper, that we are dealing with a closed system, i.e., a system completely isolated from all external influences. This idealization is the starting point of traditional thermodynamics [tHW66], according to which the second law of thermodynamics, in the formulation of Clausius ([tHW66], p.66) may be stated:
Second law (Clausius) Spontaneous changes in an adiabatically closed system will always be in the direction of increasing entropy, which rises to a maximum.
We shall refer to the above statement as the dynamical formulation of the second law, assuming, as it happens in Nature, that it refers to “spontaneous changes in time”. Given that the von Neumann entropy is a functional of the time-evolved state, we may express the dynamical formulation of the second law as
$S_{\Lambda}(\omega_{\Lambda,t_{1}})\leq S_{\Lambda}(\omega_{\Lambda,t_{2}})\mbox{ if }t_{1}<t_{2}$ (38)
Since, however, $t_{2}<-t_{1}$,  (21) yields
$S_{\Lambda}(\omega_{\Lambda,t_{2}})\leq S_{\Lambda}(\omega_{\Lambda,t_{1}})$ (39)
Equations  (38) and  (39) yield, together:
$S_{\Lambda}(\omega_{\Lambda,t_{2}})=S_{\Lambda}(\omega_{\Lambda,t_{1}})$ (40)
Equation  (40) is the expression of Schrödinger’s paradox, see [Schr65] and Lebowitz’s inspiring review [Leb08].
The reason why we have been so explicit in the above statement is that a variant of it also applies to infinite systems, as we now explain. We have seen that the notion of state generalizes to systems with infinite number of degrees of freedom $\omega(A)=\lim_{\Lambda\nearrow\infty}\omega_{\Lambda}(A)$, at first for $A\in{\cal A}_{L}$ and then to $\cal A$}. The state $\omega_{\Lambda}^{t}(A)=\omega_{\Lambda}(\exp(iH_{\Lambda}t)A\exp(-iH_{\Lambda}t))$ does not have a limit as $\Lambda\nearrow\infty$ because the spectrum is discrete and the state an almost periodic function of $t$. The functional $S_{\Lambda}$ is a continuous functional of the state $\omega_{\Lambda,t}$, and, thus, also a quasi-periodic function of $t$. As we shall see now, for states of infinite systems, however, the situation is entirely different!
For a large system the mean entropy is the natural quantity from the physical standpoint:
$s(\omega)\equiv\lim_{\Lambda\nearrow\infty}(\frac{S_{\Lambda}}{|\Lambda|})(\omega)$ (41)
The mean entropy has the following two properties [LanRo68]:
  • $a.)$
    $0\leq s(\omega)\leq\log D$ where $D=2S+1$;
  • $b.)$
    $s$ is upper semicontinuous, that is $limsup_{n\to\infty}s(\omega_{n})\leq s(\omega)$.
In b.), $\omega_{n}$ is a sequence of states such that $\omega_{n}\to\omega$ in the weak*- topology, i.e., $\omega_{n}(A)\to\omega(A)\forall A\in{\cal A}$.
We shall exclusively use the weak*-topology in the present paper. In the uniform or norm topolgy, the mean entropy is continuous, by a theorem of Fannes [Fa73], but there is no physical requirement which would reinforce the extra condition of uniformity defining the norm topology. We may call this a condition of physical naturalness: it is very important, and, indeed, crucial. Our switching to “densities”, such as the mean entropy, is also of this nature, because these are the quantities of physical relevance- the “measurable quantities”.
One simple example of b.) is a characteristic function of a closed set. The mean entropy’s basic property is the property of affinity ([LanRo68], [BRo297]):
$s(\alpha\omega_{1}+(1-\alpha)\omega_{2})=\alpha s(\omega_{1})+(1-\alpha)s(\omega_{2})\mbox{ with }0\leq\alpha\leq 1$ (42)
We now come back to the problem of the “arrow of time”.
According to Penrose [Pe79], the following is a list of phenomena associated to the “arrow of time”: “the second law of thermodynamics, the expansion of the Universe,. the use of retarded potentials in electrodynamics, the decay of the $K^{0}$-meson, and our subjective experience that the time passes”.
Schrödinger’s paradox is also one of these phenomena, because, being a consequence of the time-reversal invariance of the dynamical laws, it can only be overcome by the existence of a time-arrow, which, as in the example of the dyadic map, may be associated to the existence of the mathematical limit of infinite time for densities, whose physical significance, explained in [Wre22], is of a universal time, much greater than the typical relaxation times of systems (in a given universality class). This is very similar to the situation in statistical mechanics, where mathematical discontinuities in the thermodynamic functions arise only in the thermodynamic limit. We thus propose the following modified version of the second law:
Modified Second law (Clausius) The mean entropy of an adiabatically closed system rises monotonically to its maximum value.
Strict uppersemicontinuity of the mean entropy, as in item b.) above requires the infinite time limit : the limit $n\to\infty$ in item b.) above is short for the limit $t_{n}\to\infty$, where $t_{n},n=1,2,\cdots$ represents any sequence of time values tending to infinity.
Proposition 2.2.
Let the general interaction potential in  (28) satisfy
$|\Phi(\{0,x\})|<C(1+|x|)^{-\alpha}\mbox{ with }\alpha>2\nu$ (43)
Then, modified second law (Clausius) does not hold, unless, either the mean entropy remains constant in time, or a time-arrow exists.
Proof.
By Theorems 2.6 and 3.1 of [Wre22], see also [NSY19], under the condition  (43), for any two fixed values of the time $t_{n_{1}}$ and $t_{n_{2}}$,
$s(\omega_{t_{n_{1}}})=s(\omega_{t_{n_{2}}})=s(\omega_{t_{0}})\equiv s_{0}$ (44)
The above equation is the precise infinite-volume analogue of  (38), which expresses Schrödinger’s paradox. If strict upper semicontinuity is assumed in item b.) above, together with time-reversal invariance, we obtain
$s(\omega_{-\infty})>s_{0}$ (45)
with the interpretation of $s(\omega_{-\infty})$ as the mean entropy at the “time-reversed” universal time, contradicting the monotonic increase of the mean entropy assumed in modified second law (Clausius). ∎
2.3 Two different ways to introduce a time arrow: 1) The Boltzmann approach; 2) Adiabatic transformations and the Lieb-Yngvason approach
We have seen in  (2.2) that replacement of the entropy of a system of finite ndof by the mean entropy (a “density”) does not per se overcome the Schrödinger paradox: a time-arrow must be assumed. As it happens, however, currently accepted physical theories (with the major exception of those describing the expansion of the Universe, see the epilogue) are time-reversal invariant. Moreover, the corresponding dynamics provide a good description of equilibrium states: KMS states for $T>0$, ground states for $T=0$, see section 2.4.1.
Although it was Boltzmann who discovered (from “nothing”) the notion of entropy, and the Gibbs-von Neumann entropy is a generalization of his $\log(\rho_{\Lambda})^{\prime\prime}$ version which matches the “microcanonical ensemble”, he was certainly keenly aware of the Schrödinger paradox (see [Leb08]), which led him to introduce a new notion of entropy, today known as the Boltzmann entropy (see [Leb08] for a very illuminating introduction to this concept). It was most clearly stated and explained by Narnhofer [N25] in her recent preprint that, actually, the Boltzmann equation is an Ansatz for a dynamical evolution, and, under reasonable hypotheses [N25], equilibrium (KMS) states are globally stable. There are three main aspects of the Boltzmann approach: 1) the Boltzmann time-evolution already contains a time-arrow, and therefore does not per se require the concept of adiabatic transformation; 2.) the Boltzmann time-evolution does not provide (and does not pretend to provide) a description of equilibrium states; and 3.) the (mean) entropy does not, in general, necessarily rise to a maximal value.
The Boltzmann entropy is closely related [PPV21] to the R.A. Fisher information, which was proved to be monotonic for the Boltzmann equation in [ISV25].
In contrast, our proof of the modified second law (Clausius) in the forthcoming section 3 relies heavily on the concept of adiabatic transformation (see section 2.4.2), which is a generalization of the barrier model, because its so-called “preparation part” introduces a time-arrow. In this sense, we are closer in spirit to the approach of Lieb and Yngvason [LY99], who rely (heavily) on their notion of “adiabatic accessibility”. For systems with finite ndof their work provides a complete, axiomatic approach to the second law, in the form of a precise version of the Kelvin-Planck statement. We attempt, in section 3, to provide a dynamic complement to their work; many of their basic intuitive ideas have been, however, incorporated into our treatment. In addition, since we do not change the dynamics, (initial) equilibria continue to be well described by the associated KMS and stability conditions (see section 2.4.1), and final states are non-equilibrium steady states (NESS), associated to maximal mean entropy and “maximal mixedness”.
2.4 Dynamical properties of states of infinite systems related to the modified second law
2.4.1 Equilibrium properties, basic structural transitions underlying the modified second law and Boltzmann’s dictum
Our choice of the initial state $\omega_{0}$ as an equilibrium state may also be expected to be expressible in terms of some properties of the dynamics, viz., of the function $t\to\omega_{t}$. For temperature $T>0$ this is the well-known KMS boundary condition, see [Hu72]. We shall restrict ourselves in section 3 to the case $T=0$, and the analogous ground state condition is also well-known ([Hu72], Theorem 3.2): there is a function $F(z)$, analytic for $Imz<0$, uniformly bounded for $Imz\leq 0$, continuous on the real axis, such that, for $z=t$ (real)
$F(t)=\omega(\tau_{t}(A)B)\forall A,B\in{\cal A}$ (46)
Let $H$ be the GNS Hamiltonian associated to the state $\omega$. We say that $\omega$ satisfies the stability condition iff
$H\geq 0$ (47)
on the associated GNS space. By ([Hu72], Theorem 3.4), a state which satisfies the ground state condition is invariant and satisfies the stability condition, and vice-versa.
An initial equilibrium state may undergo, in the limit of infinite time, a basic structural change, for instance from a pure to a mixed state, or from a mixed state to a “maximally mixed state”. Note that such transitions may occur only between states of infinite systems. We shall henceforth always assume that we have to do with an initial state $\omega_{0}$ such that $\omega_{t_{n}}\to\omega$ in the weak*-topology, where $t_{n},n=1,2,\cdots$ is any sequence of time-values tending to infinity.
Proposition 2.3.
Assume  (43). If the initial state $\omega_{0}$ is a convex combination of a fixed number of pure states, and the final state is the tracial state  (30), then modified second law (Clausius) holds in the strict sense. Furthermore, each of the finite-volume approximants of the final state is the “most mixed state” in the sense of Lemma 2.2 of [NT12]. Their limit (in the weak*-topology) is also the “most mixed state”, in the sense of being associated to maximal mean entropy. Finally, the tracial state incorporates in a natural way Boltzmann’s “maximum wealth of microstates”.
Proof.
If the initial state is a convex combination of a fixed number of pure states, its mean entropy is zero due to the property of affinity  (42), and the facts 1.) that the mean entropy of a pure state is zero and 2.) the mean entropy of the tracial state equals the maximum value $\log 2$. Facts 1.) and 2.) follow under condition  (43) by theorems 2.6 and 3.1 of [Wre22], see also [NSY19].
Lemma 2.2 of [NT12] asserts that, for a density matrix $\rho$, if one measures “mixedness” by the quantity $Tr(\rho(1-\rho))=1-D(0,\rho)$, where, for any two operators $A$ and $B$, $D(A,B)\equiv Tr((A^{*}-B^{*})(A-B))$ and $A^{*}$ denotes the Hermitian conjugate operator, then the quantity $D(0,\rho)$ is a strictly convex functional of $\rho$. Indeed, for $0\leq\alpha,\beta\leq 1$, the Schwarz inequality yields $D(\alpha\rho_{1}+\beta\rho_{2})\leq(\alpha^{2}+\alpha)Tr(\rho_{1}^{2})+(\beta^{2}+\beta)Tr(\rho_{2}^{2})$. As a consequence, the “mixedness” of a state grows with the number of states in a convex combination, and is attained at the “most mixed state”, the tracial state, which is the finite-volume version of  (30). Although the proof holds only for finite number of ndof, if one measures “mixedness” of a state $\omega$ by the value $s(\omega)$, the tracial state  (30) is, again, the “most mixed state”, because it exhibits the highest value $\log 2$ of the mean entropy. Consider, now, the set of states
$\omega_{\Lambda_{1},\Lambda_{2},\Lambda_{3}}\equiv\otimes_{x\in\Lambda_{3}}[1/2(\omega_{+,x}+\omega_{-,x})]\otimes\omega_{1,2}$
where
$\omega_{1,2}\equiv\otimes_{x\in\Lambda_{1}}\omega_{+,x}\otimes_{x\in\Lambda_{2}}\omega_{-,x}$
with $\Lambda_{1}\cup\Lambda_{2}\cup\Lambda_{3}=\Lambda$ and $\omega_{\pm,x}\equiv_{x}(\pm|,\dot{|}\pm)_{x}\forall x\in\mathbf{Z}^{\nu}$. Above, $|\pm)_{x}$ are the eigenvectors of $\sigma^{3}_{x}$, and $x$ is any vector in $\mathbf{Z}^{\nu}$. Consider the limit, if it exists in the weak*-topology, of
$\lim_{\Lambda\nearrow\mathbf{Z}^{\nu}}\omega_{\Lambda_{1},\Lambda_{2},\Lambda_{3}}=\tilde{\omega}$
We have seen that the “most probable state” is $\tilde{\omega}=\omega_{tr}$, the tracial state, if we “measure” the probability by the mean entropy. If we call “microstates” the states $|\pm)_{x}$, the “most probable state” corresponds to the choices $|\Lambda_{3}|=O(|\Lambda|)$, together with $|\Lambda_{1}|=o(|\Lambda|$ and $|\Lambda_{2}|=o(|\Lambda|$. This may be proved under the condition  (43) and is a precise formulation of Boltzmann’s dictum that the “most probable state” corresponds to the “maximum wealth of microstates”. ∎
The above proposition shows that basic structural changes such as  $\mbox{ pure }\to\mbox{ mixed }$, or  mixed  most $\mbox{ mixed }\to\mbox${ most mixed } are the cornerstones of the modified second law (Clausius). On the other hand, by  (2.2), the dynamics must also allow a “time-arrow”. This leads us to our next subject.
2.4.2 Adiabatic transformations as generalizations of the barrier model. Sudden interactions
Concerning irreversible processes, Lieb and Yngvason assert in their basic paper [LY99] (p.35): “the existence of many such processes lies at the heart of thermodynamics. If they did not exist, it means that nothing is forbidden, and hence there would be no second law”.
In this connection, an immediate question arises (see also [LY99], Theorem 2.9, “Caratheodory’s principle and irreversible processes). Caratheodory’s principle asserts that “Any thermodynamical state of a system has neighborhoods which cannot be reached by adiabatic processes” ([tHW66], p.33). This principle seems, thus, to assert that, for a process or transformation, the implication adiabatic $\Rightarrow$ reversible is true, or, equivalently, irreversible $\Rightarrow$ non-adiabatic.
But, in connection with the last implication, what about the free irreversible adiabatic expansion of a gas ([tHW66], 2.4.2, p.27)?
Switching for a moment to conventional thermodynamics, with $dQ$ denoting the “heat exchange” (a concept criticized by Lieb and Yngvason for good reason), if $dQ=0$ is posed as a “definition” of adiabatic transformation, the formula
$\frac{dQ}{T}=dS$ (48)
“defining” the entropy $S$ seems to imply that $dS=0$ (reversibility) if $T\neq 0$ throughout. But, in  (48), $Q$ is not a function of state (it depends on the process). The same happens to $T$, which, in  (48) is not a number, but a function which depends on the process- as is very well-known! Indeed, a counterexample to the implication irreversible $\Rightarrow$ non-adiabatic is provided by the following
Remark 2.1.
The barrier model We consider $N$ particles of an ideal gas, initially inside a container of volume $V$, which is allowed to expand adiabatically into another container of the same volume $V$ (for simplicity), in which, originally, vacuum had been established (Gay-Lussac experiment, [tHW66], p.36). The change
$\Delta S\equiv S(U,2V,N)-S(U,V,N)$
of the entropy $S$, with the internal energy $U$ fixed (it is a constant of the motion) is found to be [tHW66]
$\Delta S=kN\log(2)\mbox{ or, equivalently }\Delta s=k\log(2)$
In order to obtain the above formula, it was assumed that the gas will fill the second container uniformly, when equilibrium is attained. Although not mentioned in [tHW66], this assumption depends crucially on two points: a.) after removal of the barrier, a new equilibrium state is established for the system; b.) this state must be compatible with the ab-initio assumptions on the states of the theory, corresponding to a given universality class. In the present case, translation-invariance  (18) determines the final configuration. Since we are dealing with a finite system, this assumption is replaced by that of periodic boundary conditions, as mentioned after  (18).
A generalization of the barrier model leads to the concept of adiabatic transformations. We have seen that they may be irreversible, in agreement with second law (Clausius), but it remains to find a proper generalization of the concept, which is both physically natural and introduces a mathematically precise concept of “sudden perturbations”. Somewhat surprisingly, the physics literature on such perturbations seems to be very scarce, an exception being Pauli’s text [Pa33]. For systems with finite number of ndof, Lieb and Yngvason solved this problem by introducing the alternative concept of “adiabatic accessibility” [LY99]. One further aspect is important in our generalization of the barrier model: the initial state. It is an equilibrium state in our tratment of the barrier model, which evolves into a non-equilibrium state by the “preparation part” of the adiabatic transformation: this is the “new” equilibrium state in part a.) above, after removal of the barrier. In the original barrier model, this new non-equilibrium state instantaneously evolves into the final state, which occupies both containers uniformly, as being the only configuration satisfying the ab-initio conditions imposed on the states. Our generalization removes the instantaneous transition, allowing for a class of perturbations taking place within a non-zero time interval.
A formulation of the concept of adiabatic transformation along the lines we present was first done by Penrose [Pe79], but without the concept of sudden interactions.
Definition 2.4.
Let a C*-dynamical system $({\cal A},\tau_{t})$ be given, with $t\in[-r,\infty)$. An adiabatic transformation of a given (initial) state $\omega_{0}$ consists of three successive steps. The first and third steps are dynamical automorphic evolutions of the state $\omega_{0}$ for t<rt<-r and $t>0$, of the form
$t\in\mathbf{R}\to\omega_{t}\equiv\omega_{0}\circ\tau_{t}$ (49)
where the circle denotes composition, i.e., $(\omega\circ\tau_{t})(A)=\omega(\tau_{t}(A))$, and $\tau$ denotes a given automorphism group under which $\omega_{0}$ is invariant. The second step, from $r\leq t\leq 0$, called preparation of the state, starts at some t=rt=-r, with $r>0$, when the state $\omega_{-r}$ is invariant under the automorphism $\tau_{-r}$, and ends at $t=0$, when $\omega_{0}$ is invariant under the automorphism $\tau_{0}$; we require the cyclicity condition
$\tau_{-r}=\tau_{0}=\tau^{1}$ (50)
It is assumed to be nontrivial, i.e., the state after preparation is not invariant under the evolution:
$(\omega_{0}\circ\tau^{1})\circ\tau_{t}\neq\omega_{0}\circ\tau^{1}\mbox{ for }t>0$ (51)
Remark 2.2.
Sudden perturbations If
$\tau^{1}=\tau$ (52)
the automorphism is continuous at t=rt=-r and $t=0$. The case
$\tau^{1}\neq\tau$ (53)
is a mathematically precise definition of a sudden interaction: it represents a “finite discontinuity” which replaces the (mathematically) ill-defined $\delta(t)$- infinity” of the lifting of barriers in constrained equilibria ([Pa33]), such as the barrier model of remark  2.1.
Remark 2.3.
Note that the automorphism $\tau$ in definition  2.4 refers to a time-independent interaction. Between t=rt=-r ant $t=0$ in the above definition a “time-dependent Hamiltonian” is supposed to act. Indeed, according to Definition  2.4, the system is closed from $t=0$ to any $t>0$, but not from t=rt=-r to $t=0$, where it is subject to external conditions, but is still thermally isolated. As informally observed in ([LY99], p.23): “Hence, as far as the system is concerned, all the interaction it has with the external world during an adiabatic process can be thought of as being accomplished by means of some mechanical or electrical devices”. This remark by Lieb and Yngvason agrees with the intuition which guided our definition  2.4: the “preparation part” of the adiabatic transformation thus defined represents an interaction with the environment of the form preconised by them. The case of sudden perturbations  (53) also illustrates their remarks on p.17: “The word ’adiabatic’ is sometimes used to mean ’slow’ or ’quasi-static’, but nothing of the sort is meant here. Indeed,an adiabatic process can be quite violent. The explosion of a bomb in a closed container is an adiabatic process”. Sudden interactions play an important role in our proof of the modified second law in section 3.
One last, but very important property of an adiabatic transformation is
Theorem 2.5.
time-arrow theorem The preparation part of a system undergoing an adiabatic transformation as defined in  2.4 necessarily breaks time-reversal invariance, that is, there exists some scalar function of the states and observables of the system which is not invariant under time-reversal.
Proof.
By  (51), there exists some observable $\tilde{A}\in{\cal A}$ such that the function
[…]
is such that
$f_{\omega_{0}}(t,\tilde{A})\neq 0\mbox{ for }t>0$ (55)
but, by invariance of $\omega_{0}$ under $\tau_{0}$, $\forall A\in{\cal A}$,
$f_{\omega_{0}}(0,\tilde{A})=0$ (56)
Thus, there exists $\tilde{A}\in{\cal A}$ such that the graph of the function $t\to f_{\omega_{0}}(0,\tilde{A}$ is not flat, and is, therefore, manifestly not time-reversal invariant. Therefore, the preparation of the system undergoing an adiabatic transformation as defined in  2.4 breaks time-reversal invariance. ∎
2.4.3 Quantum chaotic states: the role of dynamics and the classical character
The analog of “chaotic state” in section 1 for quantum systems may hold, rather surprisingly, for a class of examples of qss. Take, for instance, in the weak*-topology
$\omega_{1,2}^{\pm}=\lim_{\Lambda\nearrow\mathbf{Z}}\omega_{\Lambda,1,2}^{\pm}$ (57)
where
$\omega_{\Lambda,1,2}^{\pm}\equiv\otimes_{x\in\Lambda}|\pm)_{x,1,2}\mbox{ with }\sigma_{x}^{1,2}|\pm)_{x,1,2}=\pm|\pm)_{x,1,2}$ (58)
Then, by ([Ra70], (6), p.2950), for the Ising model in dimension $\nu=1$,
$\omega_{1,2}^{\pm}(\tau_{t}(\sigma_{x}^{1,2})=\prod_{j\in\mathbf{N}}\cos^{2}(2\epsilon(|j|)t)$ (59)
For the Dyson model,
$\epsilon(|j|)=\frac{1}{(|j|)^{s}}\mbox{ with }j\neq 0,s>1$ (60)
we find
$\omega_{1,2}^{\pm}(\tau_{t}(\sigma_{x}^{1,2})=Cl_{1/2;s>1}^{2}$ (61)
where $Cl$ denotes the Cloitre function introduced by Albert and Kiessling [AK17], given by
$Cl_{p;s}(t)\equiv\prod_{n\in\mathbf{N}}(1-p+p\cos(n^{-s}t))$ (62)
The strong chaotic fluctuations of $Cl_{p;s}(t)$ for large $t$ are given in the figure below, which is Fig.2 of [AK17], here reproduced with the kind permission of Michael Kiessling:
It may be expected from the figure that, for times $t=1,\cdots,n$, exponential sensitivity to initial conditions holds, i.e.,
[…]
The above means that any initial imprecision in the knowledge of initial conditions becomes exponentially amplified in the course of time. We are far, however, from being able to prove this conjecture! The above represents, nevertheless, strong graphical evidence that quantum chaos may exist for states of infinite systems!
It is important to observe that quantum chaos is restricted to a given universality class of interacting systems. The exponential model  (25) has trivial, soluble dynamics, with a decay of the function $\omega_{1,2}^{\pm}(\tau_{t}(\sigma_{x}^{1,2})=\sin(t)/t$ [Em66], while the corresponding functions in the Dyson models display the behavior depicted in the present section. Given that $\nu=1$, the exponential model displays no phase transition, while the Dyson models do (as first shown by Dyson in [Dy69]). Therefore, the mechanism of approach to equilibrium depends on the universality class, see the epilogue.
For $\nu=1$ and only nearest-neighbor interactions, the Ising model is equivalent to a Fermi lattice gas: the interaction contains a term
$H_{\Lambda,I}\equiv-1/2\sum_{x\in\Lambda}a_{x}^{*}a_{x+1}^{*}a_{x+1}a_{x}$ (63)
where $a_{x},a_{x}^{*}$ are operators satisfying mixed commutation and anticommutation relations $a_{x},a_{x}^{*}]=0$ if $x\neq y$, but ${a_{x},a_{x}^{*}}=1$ ([BRo297], p.426). In contrast, the “XY part” is a quadratic form in the creation-annihilation operators.
The chaotic nature of the dynamics does vary strongly with certain parameters, particularly with the time. The reader should compare Fig.2 of the paper by Albert and Kiessling, reproduced in the present paper, with their Fig.1, for small times, which displays significantly less chaos. We may understand this phenomenon intuitively if we recall that chaotic dynamics is a generic aspect of classical mechanics (classical dynamical systems), and we expect that the classical limit is defined by a condition of the form
$\frac{\hbar}{ET}\ll 1$ (64)
where $E$ denotes an energy value in the typical range allowed by the system, and $T$ a corresponding typical time value: the quantity $\frac{\hbar}{ET}$ is of course dimensionless. In our case, we may take in the Dyson model $E$ as the maximum value in  (60) and $T$ as the time (see  (59)). In this sense, comparison between the two figures in [AK17] also provides strong graphical evidence of the dictum: the limit of large times is classical!
3 The second law as a theorem for two classes of models of interacting spins at $T=0$: the necessity of extended perturbations
In this section we prove the modified second law (Clausius) for two classes of models. In order to do so, we must add some definitions.
A factor or primary state is defined by the condition that the center
$Z_{\omega}=\{\lambda\mathbf{1}\}$ (65)
with $\lambda\in\mathbf{C}$.
For quantum spin systems the center $Z_{\omega_{\beta}}$ coincides with the so called algebra at infinity $\zeta_{\omega}^{\perp}$, which corresponds to operations which can be made outside any bounded set. As a typical example of an observable in $\zeta_{\omega}^{\perp}$, let $\omega$ be any translation invariant state. Then the space average of A
$\eta_{\omega}(A)\equiv s-lim_{\Lambda\nearrow\infty}\frac{1}{|\Lambda|}\sum_{x\in\Lambda}\Pi_{\omega}(\tau_{x}(A))$ (66)
exists, and, if $\omega$ is ergodic, then
$\eta_{\omega}(A)=\omega(A)\mathbf{1}$ (67)
This corresponds to “freezing” the observables at infinity to their expectation values. The following definition is abstracted from [He72], before his Lemma 1.
Definition 3.1.
Two states $\omega_{1}$ and $\omega_{2}$ are disjoint if no subrepresentation of of $\Pi_{\omega_{1}}$ is unitarily equivalent to any subrepresentation of $\Pi_{\omega_{2}}$ - otherwise they are called coherent. Two states which induce disjoint representations are said to be disjoint.
For finite-dimensional matrix algebras (with trivial center) all representations are coherent, and factor representations as well.
We have ([He72], Lemma 6): Let $\omega_{1}$ and $\omega_{2}$ be extremal invariant (ergodic) states with respect to space translations. If, for some $A\in{\cal A}$,
$\eta_{\omega_{1}}(A)=a_{1}\mbox{ and }\eta_{\omega_{2}}(A)=a_{2}\mbox{ with }a_{1}\neq a_{2}$ (68)
then $\omega_{1}$ and $\omega_{2}$ are disjoint.
Definition 3.2.
We say that a perturbation $V$ is extended iff it is a functional of the elements of all finite parts of $\mathbf{Z}^{\nu}$. It is termed local iff it is of the form $V=a(\Lambda_{0})$, where $a(\Lambda_{0})\subset{\cal A}_{\Lambda_{0}}$, with $\Lambda_{0}$ a fixed subset of $\mathbf{Z}^{\nu}$.
Theorem 3.3.
Let
$H_{\Lambda}=\frac{1}{2}\sum_{x,y\in\Lambda}J(|x-y|)\sigma_{x}^{3}\sigma_{y}^{3}\mbox{ with }J(|x|)<0$ (69)
denote the generalized ferromagnetic Ising model (gIm), and $\tau$ the corresponding automorphism corresponding to the infinite system. Then, for both the exponential model  (25), and the Dyson model  (26) with $\alpha>2$, there exists an infinite number of distinct adiabatic transformations satisfying  (53), i.e., associated to sudden perturbations, each of extended type, such that modified second law (Clausius) for $T=0$ holds. For perturbations of local type of the ferromagnetic anisotropic Heisenberg model  (23), the modified second law (Clausius) does not hold.
Proof.
Let $\tau$ denote the time-translation automorphism associated to the finite-region Hamiltonian  (69). In the step of preparation of the state in definition  2.4 of adiabatic transformation, we take
$\tau_{-r}=\tau_{0}=\tau_{1}\neq\tau$ (70)
That is to say,  (53) is satisfied, characterizing a sudden interaction, where $\tau_{1}$ denotes the automorphism defined by
$\tau_{1}\equiv\tau\circ\tau_{P}=\tau_{P}\circ\tau$ (71)
and $\tau_{P}$ is the automorphism associated to a perturbation of the original system, associated to a finite-volume operator $V_{\Lambda}$ which is assumed to commute with the finite-region Hamiltonian  (20), i.e.,
$V_{\Lambda},H_{\Lambda}]=0\forall\Lambda\subset\mathbf{Z}^{\nu}$ (72)
Above, then, $\tau_{P}$ is defined by
$\tau_{P,t}(A)\equiv\lim_{\Lambda\nearrow\mathbf{Z}^{\nu}}(\exp(itV_{\Lambda})A\exp(-itV_{\Lambda}))\forall A\in{\cal A}_{\Lambda}$ (73)
and then by extension as usual. Take $\omega=\omega_{0}^{+}$ as one of the ground states of the Hamiltonian defined by  (69), the other case, $\omega_{0}^{-}$, as well as any convex combination of the two, being similar. Choose, in  (73), $V_{\Lambda}$ as a tranverse, sufficiently strong magnetic field (which commutes with $H_{\Lambda}$):
$V_{\Lambda}\equiv\sum_{k=-\Lambda}^{\Lambda}V_{k}$ (74)
where
$V_{k}\equiv 2(-\sum_{x\in A_{k}}h_{x}\sigma_{x}^{1}+\sum_{x\in B_{k}}h_{x}\sigma_{x}^{1})$ (75)
with $h_{x};x\in\mathbf{Z}$ a sequence of positive numbers satisfying the condition
$h_{x}>\sup_{x\in\mathbf{Z};x\neq 0}(\psi^{-|x|},|x|^{-\alpha})$ (76)
where $\alpha>2$ and $\psi>1$. Above,
$A_{k}\equiv[kC+1,\alpha_{k}]\mbox{ and }B_{k}\equiv[\alpha_{k}+1,(k+1)C]$ (77)
where
$\alpha_{k}\equiv\frac{C}{2}(1+f)\mbox{ and }\beta_{k}\equiv\frac{C}{2}(1-f)$ (78)
In  (77),  (78),
$0<f\equiv\frac{A}{B}<1$ (79)
and $C$ is chosen such that
$C\geq 2B$ (80)
If $A$ and $B$ are chosen as positive integers, by  (79) $f$ ranges over all rational numbers between zero and one. With our choice $T=0$, the restriction of the initial ground state to a region $\Lambda$, $\omega_{\Lambda,0}^{+}$, shifts under conditions  (74) to  (80), to the ground state of $V_{\Lambda}$, which is the vector state $(\Psi_{\Lambda}^{f},\cdot\Psi_{\Lambda}^{f})$, with
$\Psi_{\Lambda}^{f}=\otimes_{k=-\Lambda}^{\Lambda}\Psi_{k}^{1}\otimes\Psi_{k}^{2}$ (81)
where
$\Psi_{k}^{1}\equiv\otimes_{x\in A_{k}}|+)_{x}^{1}\otimes\cdots|+)_{x}^{1}$
and
$\Psi_{k}^{2}\equiv\otimes_{x\in B_{k}}|-)_{x}^{1}\otimes\cdots|-)_{x}^{1}$
Above, by definition, $\sigma_{x}^{1}|\pm)=\pm\frac{1}{2}|\pm)$, and the sets $A_{k}$ and $B_{k}$ have been defined in  (77),  (78) and  (79), and are thus seen to satisfy $A_{k}\cup B_{k}=[kC+1,(k+1)C]$. Moreover, by  (79) and  (80), and the fact that $A$ and $B$ are positive integers, it follows that $C>1$. The limiting state
$\omega_{f}^{1}\equiv\omega_{0}^{+}\circ\tau_{1}\equiv\lim_{\Lambda\nearrow\mathbf{Z}}(\Psi_{\Lambda}^{f},.\Psi_{\Lambda}^{f})$ (82)
clearly satisfies condition  (51), i.e., it is not a convex combination of the two ground states $\omega_{0}^{\pm}$. In fact, $\omega_{f}^{1}$ is a pure state, characterized by a rational number $f$ satisfying  (79). Any two of these pure states are disjoint by  (66) and Hepp’s lemma  (68), with the analogue of $\eta_{\omega}$ in  (68) defined by
$\eta_{\omega}(\sigma^{1})\equiv s-lim_{\Lambda\nearrow\mathbf{Z}}\sum_{x\in\Lambda}\frac{\sigma_{x}^{1}}{|\Lambda|}$ (83)
The values of the “macroscopic pointer positions” are, in the present case, the different rationals satisfying  (79). We have thus a countable infinity of distinct pure states (being disjoint, they cannot be unitarily equivalent, and thus not equal, two by two), parametrized by all rationals between zero and one. They satisfy the assumption of proposition  2.3 concerning the limit state being the tracial state, by theorem  2.1. The modified second law (Clausius) then follows from proposition  2.3. Finally, the second assertion of the theorem follows from [BKR78], because the anisotropic ferromagnetic Heisenberg Hamiltonian  (23) has a gap. ∎
4 Epilogue: Why does the mean entropy increase? The Universe, the time-arrow and the role of quantum fields, and some problems posed by AQFT
4.1 Why does the mean entropy increase? The role of dynamics and universality classes in selecting the physical mechanism of approach to a NESS
We have seen that, when the mean Gibbs-von Neumann entropy is regarded as a functional of the states of infinite systems, it may increase with time under an adiabatic transformation to a maximum value. The non-existence of a time-arrow for finite systems (Schrödinger’s paradox, see [Leb08]) remains true for states of infinite systems (proposition  2.2).
There are at least two ways to introduce a time-arrow, reviewed in section  (18): through the Boltzmann equation and evolution, or through the preparation part of the adiabatic transformations of section 2.4.2, which improves on the versions of [Wre22] and [Pe79] through the inclusion of sudden interactions.
Theorem  3.3 treats two different classes of ferromagnetic models, the exponential  (25) and the Dyson model  (26), which lie in different universality classes associated to the theory of phase transitions and the critical point in statistical mechanics: given a value of the dimension (one) and a type of symmetry (discrete), they differ in the type of decay of the interaction potential: exponential or polynomial. In both cases, for $T=0$, there exists a basic structural transition between states (from pure to mixed) in the mathematical limit of infinite time (which physically represents an universal time largely exceeding a relaxation time, see also [Wre22]). Similarly, in both cases the initial state may be any convex combination of a fixed number of pure states, and the final state must be rather special, viz. the tracial state, of maximum mean entropy, which comprises with Boltzmann’s dictum of the “maximal wealth of microstates” (see proposition 2.3).
There exists, however, a profound difference in the mechanisms of approach to the final NESS: in the exponential model the dynamics is exactly soluble (with a decay of suitable expectation values of type $\frac{\sin t}{t}$ [Em66], while, in the Dyson case, it is not exactly soluble and there exists strong graphical evidence that it displays the phenomenon of quantum chaos (section 2.4.3), as a consequence of the results of Albert and Kiessling [AK17] on the Cloitre function for large arguments (time-values). We propose to associate this phenomenon to Krylov’s mechanism [Pe79] of successive “defocalizing shocks” beginning at the position of the removed barrier. Since this mechanism is believed to be the physical mechanism of approach to equilibrium, we are confident that the present framework may be, in certain cases, physically relevant. In this connection, we should mention a fourth important aspect characterizing universality classes: interacting and non-interacting systems (see, in this respect, [MRW24]). We have seen that the Ising model may be interpreted as an interacting Fermi lattice gas (see  (63)). The XY model, on the other hand, is a quadratic form in the fermion creation and annihilation operators in this same picture, and may, in this sense, be called “non-interacting”. Although the XY-model dynamics is not chaotic, the corresponding process of “return to equilibrium” is rather subtle, and a complete analysis of it may be found in [HuRo86].
5 The Universe and the time-arrow, and the role of quantum fields
With our definition of the adiabatic transformation in section 2.4.2, we succeeded in decoupling the validity of the second law from the nature of the initial state of the Universe. Boltzmann, however, thought that a reference to the initial state was inevitable (see [Leb08] and references given there). His conjecture was, however, ultimately correct, because the problem arises as soon as we attempt to apply proposition 2.3 to the Universe. The first natural question would be: why do so many adiabatic transformations occur in the Universe? Since, however, adiabatic transformations break time-reversal invariance (theorem 2.3), Boltzmann’s previous conclusion suggests itself naturally. And, indeed, the simplest model of the Universe - the “hot big-bang theory” ( [Lu99])- confirms it, because it leads to an ever-expanding Universe, which is one of the phenomena associated to the “arrow of time”. The details of the expansion depend, however, on much more refined models. One of them, due to Maeder [Ma17], explores an additional symmetry - the scale-invariance of the macroscopic empty spaces - which does not require the assumptions of “dark energy” or “dark matter”.
A preliminary model of the CMB (“cosmic background radiation”) which fills the Universe today with a black-body spectrum of temperature $T\approx 2,7K$ assumes a system of free photons, i.e., which do not interact with the Robertson-Walker metric [Lu99]. It requires only free quantum fields. In general, interacting quantum fields - in principle, described by algebraic quantum field theory (AQFT) [CH01] - will be required. An important difference between our ferromagnetic models and AQFT is the fact that, in contrast to AQFT, the separating character of local algebras does not hold, because our ground states posess annihilators among the local algebras ${\cal A}_{\Lambda}$, viz.
$\sum_{x\in\Lambda}(\sigma_{x}^{3}\pm 1/2)$
For this reason, the deep theorem in [CH01] that the set of entangled states is generic, is not necessarily true for infinite quantum ferromagnets.
6 Some problems posed by AQFT
It would be very nice to extend the present framework to AQFT, but there are several problems in this connection. Firstly, one may conjecture that physical theories such as quantum electrodynamics may be ultimately described by AQFT, but, unfortunately, as observed by Lieb and Loss [LL00], the quantity $E_{0}(\Lambda,\Gamma)$, where $\Lambda$ denotes a space cutoff and $\Gamma$ an ultraviolet cutoff, which is the ground state energy to be subtracted from the Hamiltonian $H_{\Lambda,\Gamma}$, to account for the condition of stability  (47), is not correctly described by perturbation theory, which, in the mathematically rigorous approach of “formal series”, predicts $O(|\Gamma|^{2})$, contradicting the Lieb-Loss inequalities (see [LL00] and references given there). This means that any progress depends on nonperturbative exact or rigorous results on the dynamics, which are not yet available.
Acknowledgements
I should like to thank Elliott Lieb and Heide Narnhofer, who have strongly influenced the present paper, through various scientific exchanges, very heartily. Heide also helped me to improve the final version with several important remarks. I am also indebted to the late Derek Robinson, who insisted on the relevance of the property of affinity, which plays an important role in this paper. Finally, I should like to thank Michael Kiessling for his kind permission to include Fig.2 of his paper with Albert, which turned out, unexpectedly, to describe the chaotic quantum dynamics of the whole class of Dyson models.